OpenStudy (anonymous):

another series question -- my answer disagrees with the ans in my solutions manual, and I agree with their solution but can't find the fault in mine so I'm here to look for someone to pick at my mistakes... =) will take time to throw down my sol'n

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 2\times4\times6...\times(2n) }{ n! }\]

OpenStudy (anonymous):

and I'm asked if it converges absolutely, conditionally or diverges, so I thought I'd just use the ratio test

OpenStudy (anonymous):

but after working through it all I end up with 1/n

OpenStudy (rational):

are you really sure the denominator is n! and not n!^2 ?

OpenStudy (anonymous):

\[|\frac{ 2\times4\times6...\times(2(n+1)) }{ (n+1)! } \times \frac{ n! }{ 2\times4\times6...\times(2n) }|\] and then I simplified... yeah, it's definitely just n!

OpenStudy (anonymous):

so when simplifying, is it cool to just cancel all of the 2 times 4 times 6 and keep just the 2(n+1) and 2(n)? That may be where I'm slipping up

OpenStudy (rational):

that simplifies to 2(n+1)/(n+1) which is same as 2 right ?

OpenStudy (anonymous):

I simplified to 1/n =o I'll try a third time, then, it's probably just me being a dunce

OpenStudy (rational):

first simplify the factorial terms maybe : n!/(n+1)! = 1/(n+1)

OpenStudy (anonymous):

\[|\frac{ (2(n+1)) }{ (n+1)! } \times \frac{ n! }{ (2n) }|\]\[|\frac{ (n+1) }{ (n+1)! } \times \frac{ n! }{ (n) }|\]\[\frac{ (n+1) }{ (n+1)(n!) } \times \frac{ n! }{ (n) }\]

OpenStudy (rational):

no wait, there is a mistake in first line

OpenStudy (anonymous):

am I messing with the factorials wrong? Totally new to them, also a little slow, so it's entirely possible lol

OpenStudy (rational):

that (2n) should not be there in the denominator

OpenStudy (anonymous):

But when you cancel all the product of even integers, doesn't the 2n remain? D=

OpenStudy (rational):

let me show u step by step. start with \[\left|\frac{ 2\times4\times6...\times(2(n+1)) }{ (n+1)! } \times \frac{ n! }{ 2\times4\times6...\times(2n) }\right| \]

OpenStudy (rational):

same as \[\left|\frac{ 2\times4\times6...\times(2n)\times (2(n+1)) }{ (n+1)! } \times \frac{ n! }{ 2\times4\times6...\times(2n) }\right| \] yes ?

OpenStudy (anonymous):

ahh, okay so then there's also a 2n if we're going up until 2(n+1)

OpenStudy (rational):

thats it!

OpenStudy (anonymous):

Thank you!!!!!!!!!! You've helped me a bunch tonight I hope I can repay the favour sometime

OpenStudy (rational):

lol that reminds me of godfather movie somehow